OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6?

OpenStudy (aripotta):

what do we know about parallel lines? what do they have in common?

OpenStudy (anonymous):

they never cross?

OpenStudy (aripotta):

yes. so that means they have the same...?

OpenStudy (anonymous):

slope?

OpenStudy (aripotta):

yes :D so now we can put the information into point-slope form...what would that look like?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

i have 1 minute before i have to go lol

OpenStudy (aripotta):

so once we put the info in, it'll be y - 4 = -1(x - 3) now you can distribute the -1 into the parentheses and change it into standard form

OpenStudy (anonymous):

so x+y=7

OpenStudy (aripotta):

yup :D good job